OpenStudy (woozysuace73737):

A rectangular park is 150 yards on one side and 125 yards on the other. If Debbie walks around the park two times how far does she walk. If Debbie wanted to walk 1,000,000 yards how many times she would have to walk around the park

OpenStudy (vuriffy):

Hello.

OpenStudy (vuriffy):

Do you know how to find the area of a rectangle?

OpenStudy (woozysuace73737):

No

OpenStudy (vuriffy):

Area = base * height

OpenStudy (woozysuace73737):

Ok

OpenStudy (vuriffy):

One side is 150, another side is 125. You may multiply these values to receive the area of the rectangle, do you mind doing that for me?

OpenStudy (woozysuace73737):

I can do that

OpenStudy (vuriffy):

Type your answer when you are done, thank you.

OpenStudy (woozysuace73737):

18,750

OpenStudy (vuriffy):

Yes, correct.

OpenStudy (woozysuace73737):

37,500

OpenStudy (vuriffy):

I apologize, I made a mistake.

OpenStudy (woozysuace73737):

Ok

OpenStudy (woozysuace73737):

Ok

OpenStudy (vuriffy):

How many sides are there on a rectangle?

OpenStudy (woozysuace73737):

4

OpenStudy (vuriffy):

Two sides have the same value, so 2(150) + 2(125) = ?

OpenStudy (vuriffy):

300 + 250 = one whole perimeter around the park.

OpenStudy (vuriffy):

Debbie walks around it twice, which means you multiply the perimeter by 2. 550*2 = ?

OpenStudy (woozysuace73737):

279?

OpenStudy (vuriffy):

I'm sorry where did you get that value?

OpenStudy (vuriffy):

There is two sides with the same value, so 150 + 150 + 125 + 125 = ?

OpenStudy (woozysuace73737):

My caculator

OpenStudy (vuriffy):

The perimeter is 550 yards.

OpenStudy (woozysuace73737):

550

OpenStudy (vuriffy):

Multiply by two for two times around the park. 550 * 2 = ?

OpenStudy (woozysuace73737):

1,100

OpenStudy (vuriffy):

Area = 18,750 yards Perimeter = 550, Debbie walks around it two times, so 550 * 2 = 1,100. Now the final part: All you have to do is 1,000,000 yards divided by the perimeter, which is 550.

OpenStudy (vuriffy):

1,000,000 / 550 = ?

OpenStudy (woozysuace73737):

Ok

OpenStudy (woozysuace73737):

1818

OpenStudy (vuriffy):

Okay good, but she wants to go at least 1,000,000 yards.

OpenStudy (vuriffy):

So, you must add 1, because if you didn't she wouldn't have gone at least 1,000,000 yards total.

OpenStudy (vuriffy):

So, the answer would be 1819 times.

OpenStudy (woozysuace73737):

Ok thank you you ar so smart :)

OpenStudy (vuriffy):

Sure thing, anytime!

OpenStudy (woozysuace73737):

Also could you help me with a other question